By how much does r - s exceed s - r ?
a (r-s)/(s-r) b 2(r-s) c -2 d 2(s-r) e -1

Answer by stanbon(75887)

By how much does r - s exceed s - r ?
a (r-s)/(s-r) b 2(r-s) c -2 d 2(s-r) e -1
--------------------------------
Your Problem:
(r-s)-(s-r) = r-s-s+r = 2r-2s = 2(r-s)
====================
Cheers,
Stan H.